OpenStudy (anonymous):

Find the real numbers a a and b such that (a+bj)^2=16-30j ;a>0

OpenStudy (anonymous):

(a + bj)^2 = a^2 +2 ab j - b^2 = 16 - 30j 2 ab = 30 a^2-b^2 = (a+b)(a-b) = 16 substitute a = 30/2b and solve ?